### v1

@@ 6.4 Setup 4: Temporal Q-Value Analysis
Results. We trained the attacker for 40 episodes (400 iterations), logging Q-values for all visited state-action pairs. Figure 5111 For the CTF environment, actions 1–18 are: local(ScanBashHistory), local(ScanExplorerRecentFiles), local(SudoAttempt), local(ExfiltrateFlag), local(CrackKeepPassX), remote(ProbeLinux), remote(ProbeWindows), remote(ProbeSQLServer), connect(HTTPS), connect(GIT), connect(SSH), connect(RDP), connect(MySQL), connect(SSH-key), connect(PING), connect(su), remote(ProbeFlagServer), and local(CrackKeepPass). For the Chain environment, actions follow this sequence: local(CrackKeepPassX), remote(ProbeLinux), local(ScanBashHistory), remote(ProbeWindows), local(CrackKeepPass), connect(SSH), connect(RDP), connect(SSH-key), connect(MySQL), connect(HTTPS), connect(GIT), connect(PING), local(ScanExplorerRecentFiles), local(SudoAttempt), connect(su). shows Q-value distributions per action at different training stages across environments, with Table LABEL:tab:abstract-actions (refer appendix) mapping action indices to semantic descriptions. In early episodes, Q-values are relatively low and undifferentiated, indicating uncertainty and limited experience. As training progresses, a preference hierarchy emerged, with one single action (highlighted in red) attaining higher mean Q-values across the states, indicating stronger expected long-term value.
The sharpness and timing of this differentiation varies by environment. For CC22, dominant early actions include local(CrackKeepPassX), connect(RDP), and connect(GIT) (actions 1, 7, and 11) by episode 20 (Figure 5c), reflecting early experimentation with lateral movement and credential reuse. By episode 35 (Figure 5d), it converges on local(CrackKeepPass) (action 5), suggesting a shift toward more reliable or higher-reward local exploitation once basic network structure is learned. For CC100, a similar pattern is observed, the agent begins with broad action exploration, including connect(SSH-key), connect(MySQL), connect(GIT), and local(ScanExplorerRecentFiles) (actions 8, 9, 11, 13) at episode 10 (Figure 5e), before narrowing its focus to local(CrackKeepPass) (action 5) by episode 35 (Fig. 5f). This likely reflects convergence on a consistently valuable tactic once credential chains are established. For CC500, in contrast to the smaller environments, the agent ultimately prefers remote(ProbeWindows) (action 4) by episode 40 (Fig. 5h), after initial attempts using actions like local(CrackKeepPassX) and connect(RDP). This suggests that in more complex, large-scale networks, reconnaissance actions offer higher long-term strategic value, providing the visibility needed to coordinate deeper exploitation.
@@
These results reflect the agent’s shift from broad exploration to targeted exploitation of high-reward paths. In smaller environments (CC22, CC100), the agent converges on local exploits like local(CrackKeepPass) (action 5), while in larger networks like CC500, it favours reconnaissance actions such as remote(ProbeWindows) (action 4), highlighting a preference for scalable reconnaissance over direct exploitation. By tracking how action valuations evolve over time, our framework helps developers debug brittle convergence, uncover underused actions, and evaluate policy stability. For cybersecurity teams involved in red/blue-team exercises or AI-driven threat simulation, these insights offer a window into how autonomous agents learn and which tactics they prioritise, informing training scenarios, service hardening, and timing of decoy deployment during simulations.
6.5 Setup 5: Strategic Decision Attribution via PER
@@
This experiment demonstrates how Prioritised Experience Replay surfaces the most impactful (high-TD-error) transitions, acting as an attention mechanism that steers the agent toward strategically important experiences.
@@
Figure 6: Learning Impact of High-Priority Transitions (Average TD Error) Across Environments.
Results. PER was configured with $\alpha=0.6$, $\beta=0.7$, and $\epsilon_{0}=0.01$ (to avoid zero probabilities). We filtered transitions with rewards above 2.0 to isolate high-impact experiences such as privilege escalation. Figure 6 shows average transition priority over time. All environments exhibit early TD-error spikes from policy instability and exploration. Smaller networks (CTF, CC22) stabilise quickly, while larger ones (CC100, CC500) show sustained fluctuation, indicating extended learning. Notably, CC500 displays distinct spikes, suggesting newly discovered pivot points or strategic escalations. Figure 7 tracks the number of distinct high-priority states. CTF and CC22 converge to compact state sets, implying repeated exploitation of a compact, optimal state set. Larger networks (CC100, CC500) maintain broader distributions, indicating ongoing refinement and diverse learning trajectories.
PER reveals where and when attacker agents learn most rapidly during training. These TD-error spikes, especially in large, dynamic environments, mark key policy shifts that developers can investigate and defenders can simulate against. For developers, CC100, the smooth arc between 40k-60k batch iterations reflect gradual learning concentration, useful for assessing when and how the agent starts committing to high-impact transitions. For defenders, while PER is not observable in deployed agents, it supports red-team planning by revealing when attackers are most likely to pivot. In CC500, twin spikes around 15k and 25k iterations may reflect credential pivoting or lateral movement, ideal timing cues for testing decoy deployment or access hardening in simulation. Future agentic threats using online learning could exhibit similar dynamics.
@@
Figure 7: Number of Key High-Priority States Across Environments.
7 Discussion
Takeaways from Analysis. Our multi-layer XRL framework provides structured visibility into attacker behaviour at both strategic and tactical levels. At the MDP level, it surfaces transitions from reconnaissance to exploitation, contextualised by uncertainty and environment progression. At the policy level, it highlights how action preferences emerge and evolve, and which transitions drive the agent’s most significant updates. Together, these layers go beyond performance metrics to explain how and when attacker strategies are formed, enabling proactive interpretation rather than retrospective outcome assessment.
Practical Use for Cybersecurity Experts. Although designed for RL explainability, our framework offers operational benefits for security teams planning for future agentic threats. As autonomous attackers become more sophisticated, defenders must understand not only what an agent does but also how it learns and adapts. Our framework enables defenders to anticipate attacker intent by highlighting when and where adversarial strategies shift, especially during red-team development or simulation-based training.
Example: In a simulated environment like CC500, defenders might observe that a DQL agent consistently spikes in transition priority around 15k and 25k iterations, signalling key learning updates such as credential pivoting. This insight helps teams identify likely escalation paths and optimal timing for deploying decoys or adjusting access controls. While such data are not available in live attacks, it supports red-team preparation and AI adversary modelling by revealing which phases of attack progression are most adaptive.

### v2

@@ VI-D Setup 4: Temporal Q-Value Analysis
Results. We trained the attacker for 40 episodes (400 iterations), logging Q-values for all visited state-action pairs. Figure 6 shows Q-value distributions per action at different training stages across environments, with Tables II and IV defining the mapping from action indices to semantic descriptions for the Chain and CTF environments, respectively. In early episodes, Q-values are relatively low and undifferentiated, indicating limited experience and high uncertainty. As training progresses, a preference hierarchy emerges: a small number of actions become consistently favored, with one action (highlighted in red) attaining the highest mean Q-value across states, indicating stronger expected long-term value.
Learning Trajectories by Environment:
  • CTF (Figure 6(a), 6(b)): A clear pattern emerges where several actions, including ICMP ping, Attempt sudo privilege escalation, Connect to Git service, RDP connection with cached creds, and Crack KeepPassX vault (Linux) (action 3, 5, 10, 12, 15), consistently receive negative Q-values. This suggests these options are ineffective or yield poor long-term value under the current reward and exploration strategy. Moreover, the agent converges on remote(ProbeLinux) (action 6) as its highest-valued action by episode 25, indicating a deliberate pivot toward OS fingerprinting and reconnaissance. This reflects an interpretable, interpretable shift toward information-gathering tactics even in smaller environments.
  • CC22 (Figure 6(c), 6(d)): Early exploration favours local(CrackKeePassX), connect(RDP), and connect(GIT) (actions 1, 7, 11), consistent with lateral movement and credential reuse. By episode 35, the agent settles on local(CrackKeePass) (action 5), marking a transition to reliable, local credential cracking once the network is partially mapped.
  • CC100 (Figure 6(e), 6(f)): The agent begins with broad exploration, including connect(SSH-key), connect(MySQL), connect(GIT), and local(ScanExplorerRecentFiles) (actions 8, 9, 11, 13), and converges on the local(CrackKeePass) action by episode 35. This again signals the emergence of credential-based exploitation as a robust tactic once credentials become available.
  • CC500 (Figure 6(g), 6(h)): Unlike smaller environments, CC500 favours reconnaissance. The agent begins with trials of local(CrackKeePassX), remote(ProbeWindows), and connect(RDP) (action 1, 4, 7), but by episode 40, remote(ProbeWindows) (action 4) dominates. This pivot reflects a visibility-first strategy in large networks, where reconnaissance precedes coordinated exploitation. For cyber range developers/red team trainers, this validates that training regime fosters realistic, stepwise escalation rather than brittle or impulsive policies.
@@
These results highlight the agent’s progression from broad exploration to targeted exploitation of high-reward paths, with environment scale shaping the final strategy. In smaller environments (CC22, CC100), the agent converges on local exploits such as, local(CrackKeePass) while in larger networks like CC500, it prioritises reconnaissance actions such as remote(ProbeWindows) (action 4), reflecting a preference for scalable reconnaissance before deeper compromise. By tracking the evolution of Q-values over time, our framework surfaces these strategy-dependent patterns, enabling developers to diagnose brittle convergence, identify underutilised actions, and assess policy stability. For cybersecurity teams engaged in red/blue-team exercises or AI-driven threat simulation, these insights provide a clear view into how autonomous agents learn and prioritise tactics, informing training design, service hardening, and the strategic timing of decoy deployments.
VI-E Setup 5: Strategic Decision Attribution via PER
@@
This experiment demonstrates how Prioritised Experience Replay surfaces the most impactful (high-TD-error) transitions, acting as an attention mechanism that steers the agent toward strategically significant experiences.
TABLE III: Action indices and descriptions for CTF environment. These indices (1-18) correspond to the x-axis in Figure 6 and are used to interpret the emergence of action preferences.
Index Abstract Action Description
1 local(ScanBashHistory) Scan bash history for host references
2 local(ScanExplorerRecentFiles) Scan Windows recent-files list
3 local(SudoAttempt) Attempt sudo privilege escalation
4 local(ExfiltrateFlag) Read local flag file
5 local(CrackKeePassX) Crack KeepPassX vault (Linux)
6 remote(ProbeLinux) Probe to confirm Linux OS
7 remote(ProbeWindows) Probe to confirm Windows OS
8 remote(ProbeSQLServer) Probe for SQL-Server service
9 connect(HTTPS) Connect to HTTPS service
10 connect(GIT) Connect to Git service
11 connect(SSH) SSH connection with cached creds
12 connect(RDP) RDP connection with cached creds
13 connect(MySQL) Connect to MySQL service
14 connect(SSH-key) SSH using leaked key
15 connect(PING) ICMP ping
16 connect(su) Command/connection via su creds
17 remote(ProbeFlagServer) Probe the flag server
18 local(CrackKeePass) Crack KeepPass vault (Windows)
@@
Figure 7: Temporal evolution of average TD-error (PER priority) across environments. Smaller environments (CTF, CC22) stabilise quickly, whereas larger networks (CC100, CC500) exhibit prolonged surges in TD-error, signalling persistent replay of unstable transitions. These spikes correspond to the performance degradation discussed in Section VI-E (“Unexpected PER Behaviour”).
@@
Figure 8: Number of Key High-Priority States Across Environments. In smaller networks (CTF, CC22), the agent quickly converges to a small set of key states, indicating early policy convergence. In contrast, larger networks (CC100, CC500) exhibit broader, persistent distributions, reflecting prolonged exploration and slower convergence due to higher state-space complexity.
Results. We trained the agent for 20 episodes (1500 iterations each) in CTF, 20 episodes (2000 iterations each) in CC22, 100 episodes (2000 iterations each) in CC100, and 200 episodes (4000 iterations each) in CC500. PER was configured with $\alpha=0.6$, $\beta=0.7$, and $\epsilon_{0}=0.01$, following standard practice in the PER literature, where $\alpha$ controls the strength of prioritisation, $\beta$ compensates for the bias introduced by non-uniform sampling via importance sampling, and $\epsilon_{0}$ prevents zero sampling probability [21]. We filtered transitions with rewards above 2.0 to isolate high-impact experiences such as privilege escalation. Figure 7 shows the evolution of average transition priority over time. All environments exhibit early TD-error spikes from policy instability and exploration. Smaller networks (CTF, CC22) stabilise quickly, while larger ones (CC100, CC500) show sustained fluctuation, indicating slower convergence and repeated reprioritisation of experience samples. In CC500, a sequence of mid-training spikes between approximately 9K and 18K batch iterations reflects intermittent bursts of partial policy reorganisation. These surges likely correspond to successive discoveries of new pivoting or credential paths, followed by unstable replay of high-TD transitions. The irregular amplitude and recurrence of these spikes suggest that, beyond initial learning gains, PER began over-emphasising transient experiences, amplifying noise and delaying convergence. Figure 8 tracks the number of distinct high-priority states across training. CTF and CC22 converge to compact sets of key states, implying repeated exploitation of an optimal state set. In contrast, larger networks (CC100, CC500) maintain broader, persistent distributions, reflecting ongoing refinement and diverse learning trajectories in larger state spaces.
Interpretation and Utility. PER visualisations reveal when and where attacker agents learn most rapidly during training. In small networks, contraction of the high-priority state set signifies convergence toward a stable attack strategy. In larger environments, intermittent TD-error surges mark critical phases of strategic exploration, information that developers can use to assess convergence timing, reward stability, and exploration schedules. For defenders, while PER signals are not observable in deployed agents, analysing these training-time patterns helps identify the states, assets, or pathways most likely to be targeted or re-used by adaptive agents, insights that can inform the design and placement of decoys, monitoring hooks, or segmentation policies in simulated or operational settings.
Explainability as a Diagnostic Lens: Revealing PER Failure Modes. PER is widely adopted to improve sample efficiency by amplifying transitions with high TD error. While often assumed to accelerate convergence, its behaviour in large, partially observable cyber environments remains poorly understood. Using our training-time explainability framework, we identify a diagnostic failure mode of PER that is not observable from aggregate reward trajectories alone. As shown in Figure 9, DQL augmented with PER (DQL+PER) exhibits less stable learning and inferior final rewards than standard DQL in both CC100 and CC500 environments.
Reward curves show the symptom; internal signals expose the mechanism. Explainability signals reveal a replay-driven distortion in the learning process that fundamentally alters value propagation at scale. Figure 10 exposes the internal dynamics responsible for this behaviour. In CC100, DQL+PER sustains elevated TD-error for a prolonged period relative to standard DQL, indicating repeated replay of high-error but weakly informative transitions that slow value stabilisation and delay convergence. In CC500, while both agents exhibit large TD-errors during early exploration, standard DQL subsequently stabilises, whereas DQL+PER shows slower decay and persistent TD-error fluctuations across later episodes. This pattern reflects misaligned replay priorities and inconsistent value propagation as the state space expands.
(a) CC100
(b) CC500
Figure 9: Reward progression of DQL and DQL+PER in CC100 and CC500 environments. DQL achieves higher cumulative rewards and steadier learning, while DQL+PER underperforms due to its bias toward high-TD-error transitions, which oversamples transient experiences and slows convergence.
Figure 10: Average TD-error of DQL and DQL+PER in CC100 and CC500. DQL learns more stably, whereas PER introduces early instability and biased convergence, leading to inferior rewards.
Inspection of PER-priority traces and temporal Q-value dynamics further indicates that many high-TD-error transitions sampled in later training stages correspond to unstable, short-lived behaviours rather than enduring strategic improvements. Continued oversampling of such transitions after their relevance has decayed injects noise into the replay buffer, disrupts value estimation, and impedes policy stabilisation. This constitutes a negative result enabled by training-time explainability: under reward-only evaluation, PER appears to underperform without clear explanation, whereas internal diagnostics reveal that the degradation arises from replay-driven distortions rather than insufficient exploration, inadequate training duration, or poor hyperparameter choices.
These results highlight a fundamental limitation of naïvely applying PER at scale. Prioritisation schemes must be co-designed with environment complexity and learning dynamics, as their effectiveness is inherently context-dependent rather than universally guaranteed. More broadly, this analysis demonstrates how training-time explainability enables the discovery of optimisation pathologies that remain opaque under performance-centric evaluation, reinforcing the necessity of internal diagnostic signals when developing and validating autonomous cyber agents.
(a) CTF: Early Phase
(b) CTF: Late Phase
(c) CC22: Early Phase
(d) CC22: Late Phase
(e) CC100: Early Phase
(f) CC100: Late Phase
(g) CC500: Early Phase
(h) CC500: Late Phase
Figure 11: Cumulative rewards comparison for the PPO algorithm in early vs. late attack phases across environments. Results show that PPO agent achieves higher rewards in the late phase, indicating a progression from exploratory behaviour to exploitation-focused strategies as training advances.
(a) CTF: Episode 10
(b) CTF: Episode 25
(c) CC22: Episode 20
(d) CC22: Episode 35
(e) CC100: Episode 10
(f) CC100: Episode 35
(g) CC500: Episode 5
(h) CC500: Episode 40
Figure 12: Emergence of action preferences through state-aggregated policy logits across episodes. The highlighted dominant action marks the shift from exploratory to high-value, environment-specific tactics. Compared with DQL, PPO’s clipped updates yield smoother yet consistent preference consolidation. Action indices for CyberBattleChain (1–15) and CTF (1–18) correspond to the x-axis and are listed in Table II and Table IV, respectively.
VI-F Cross-Algorithm Generalisability of the Explainability Framework
While Sections VI-A to VI-E establish detailed interpretability results under Deep Q-Learning, a fundamental question remains: do these findings reflect intrinsic properties of adversarial learning, or are they artefacts of a particular optimisation strategy? Our framework is explicitly model-agnostic, designed to extract explainability signals from any RL agent that exposes either value estimates or policy outputs. To substantiate this claim, we replicate two representative experimental setups, Phase-aware behavioural evolution (Setup 3) and Temporal value dynamics (Setup 4), using the policy-gradient Proximal Policy Optimisation (PPO) algorithm with matched network architectures. This cross-paradigm evaluation probes whether the strategic and tactical explainability signatures identified earlier, such as phase transitions, action consolidation, and learning stability, remain invariant under fundamentally different optimisation mechanisms, thereby establishing the framework’s generality across learning paradigms rather than dependence on any single algorithm.
VI-F1 Strategic-Level Generalisability: Phase Transitions and Behaviour Evolution
Figures 5 (DQL) and 11 (PPO) show highly consistent phase-aware reward trajectories across both learning families and all evaluated environments. Despite their fundamentally different optimisation mechanisms, bootstrapped value updates in DQL versus clipped policy-gradient optimisation in PPO, both agents exhibit the same macro-level behavioural progression: an exploratory reconnaissance phase with sparse rewards, followed by a decisive exploitation phase marked by rapid reward growth and stable convergence. This pattern mirrors the phase structure identified for DQL in Section VI-C and persists under PPO, demonstrating that the exploration–exploitation transition is not an artefact of value-based learning but a stable structural property of adversarial training dynamics. The figures further show algorithm-specific differences superimposed on this shared structure: DQL displays sharper reward inflection points as critical vulnerabilities are discovered, whereas PPO produces smoother, more gradual transitions with narrower confidence bands, consistent with its variance-controlled updates and entropy regularisation.
The results show that the timing and qualitative form of the phase transition remain consistent across both algorithms and network scales, indicating that the strategic dynamics revealed by our framework arise from the underlying structure of adversarial learning itself, rather than from the particulars of any individual optimisation algorithm.
VI-F2 Tactical-Level Generalisability: Temporal Value Evolution and Action Preferences
At the tactical level, Figures 6 (DQL) and 12 (PPO) reveal a consistent pattern in the temporal emergence of dominant actions, a principal indicator of tactical consolidation. Across both value-based and policy-gradient agents, early training is characterised by diffuse exploratory preferences, followed by progressive concentration onto a small subset of high-impact, environment-specific actions. In CTF environment, both algorithms initially prioritise reconnaissance-oriented actions (e.g., connect(PING), connect(MySQL)), before converging on remote(ProbeLinux) as the dominant exploit, marking the transition from exploration to exploitation. In the more complex CC500 environment, both DQL and PPO ultimately favour critical pivoting actions such as remote(ProbeWindows) and connect(GIT), despite substantial differences in their underlying optimisation dynamics. While DQL’s Q-values exhibit sharper episodic fluctuations and discrete shifts in dominant actions, PPO’s policy logits evolve more smoothly with broader preference distributions, reflecting their contrasting update mechanisms: DQL’s bootstrapped value updates induce higher variance, whereas PPO’s clipped-gradient formulation promotes gradual policy refinement. Occasionally, PPO exhibited locally over-confident policy behaviour, where one action’s preference sharply dominated others late in training. Although this reflects expected policy-gradient saturation under sparse-reward conditions, it also demonstrates the framework’s ability to surface such optimisation dynamics for inspection. These parallel dominance trajectories reveal that the same subset of actions preserves relative importance across algorithms, even as value scales and convergence rates differ. Collectively, this demonstrates that the hierarchy of action importance remains structurally aligned between DQL and PPO, confirming that the framework captures algorithm-independent tactical regularities in adversarial RL.
(a) Evolution of policy confidence and uncertainty.
(b) Reward and policy entropy dynamics over training.
Figure 13: Explainable signatures of policy collapse. (a) Evolution of policy confidence (logit margin) and uncertainty (normalised entropy) over training. A sharp rise in confidence and collapse in entropy mark the transition to an over-confident, low-exploration regime. (b) Reward–entropy dynamics during training. Following the entropy collapse, reward stagnates, indicating behavioural failure despite high confidence.
Across both value-based (DQL) and policy-gradient (PPO) agents, the same strategic and tactical explainability signatures consistently emerge. Phase transitions from exploration to exploitation, consolidation of dominant actions, and stability patterns persist despite fundamentally different optimisation mechanisms. This demonstrates that the signals extracted by our framework reflect intrinsic properties of adversarial learning dynamics rather than artefacts of a particular RL algorithm, confirming the model-agnostic nature of the proposed explainability approach.
VI-G Policy-Level Explainability: Internal Confidence–Uncertainty Dynamics
To evaluate the diagnostic power of our explainability framework under the most challenging conditions, where learning appears stable under reward-based evaluation, we analyse PPO trained in the CC500 environment. CC500 represents a large-scale, more complex, and highly partially observable setting, while PPO is widely regarded as a robust, entropy-regularised optimisation method. Failure observed in this regime therefore cannot be dismissed as an artefact of value-based learning or small environments.
This experiment isolates the internal dynamics of the PPO policy in the CC500 environment to determine whether observed learning failures originate within the policy itself. We deliberately exclude task-level performance signals in order to identify structural failure mechanisms before they become observable in reward trajectories. At each training episode, we instrument two complementary policy-level signals: policy entropy, capturing decision uncertainty and exploratory capacity, and the logit margin (the difference between the highest and second-highest action logits), capturing instantaneous decision confidence. Their temporal evolution is shown in Figure 13(a). During early training (Episodes 1–18), the policy operates in a flexible regime: entropy remains high, logit margins are near zero, and action preferences are weakly differentiated. This configuration reflects sustained exploration and effective uncertainty management. Between Episodes 19–25, a sharp structural transition occurs. Policy entropy collapses rapidly toward zero while the logit margin rises abruptly, forcing the policy into a highly concentrated, near-deterministic configuration. From this point onward, the agent exhibits persistent over-confidence, vanishing uncertainty, and a near-complete loss of exploratory behaviour. Crucially, this internal policy collapse occurs before any corresponding degradation in task-level performance, indicating that failure emerges within the policy rather than as a response to declining reward.
Figure 13(b) contextualises this internal degeneration by relating entropy dynamics to reward evolution. Prior to collapse, high entropy coincides with volatile but improving reward. Once entropy contracts, reward becomes unstable and subsequently deteriorates, revealing a growing misalignment between internal decision commitment and effective behaviour. These structural breakdowns are not detectable under reward-only evaluation.
By exposing the contraction of the policy’s decision space during training, this analysis establishes the internal origin of learning failure and provides an early-warning signal that precedes performance collapse. The following section examines how this internal degeneration propagates to produce global performance failure.
Policy entropy and confidence dynamics reveal that learning failure originates inside the policy rather than in task performance. The agent enters an over-confident, low-uncertainty regime while rewards are still improving, demonstrating that policy-level explainability can surface latent failure mechanisms that reward-centric evaluation cannot detect.
Figure 14: Joint evolution of task performance (normalised return), exploration (policy entropy), and decision confidence (policy sharpness) over training.
VI-H Joint Explainability of Policy and Performance Collapse
This experimental setup examines the temporal relationship between internal policy degradation and task-level performance collapse. Figure 14 jointly visualises policy entropy, policy sharpness, and normalised return, enabling direct comparison of internal learning dynamics with external behavioural outcomes. During early training (Episodes 1–18), the agent operates in a stable learning regime: high policy entropy sustains exploration, policy sharpness remains low, and task performance improves steadily. These signals evolve coherently, indicating effective optimisation. Between Episodes 19–26, a critical transition occurs. Policy entropy contracts rapidly while policy sharpness saturates, forcing the policy into an over-confident, low-uncertainty regime. Importantly, this internal collapse precedes any degradation in task-level performance. Only after the policy has entered this degenerate state does task performance collapse and fail to recover, marking the onset of persistent learning failure. This temporal ordering establishes internal policy dynamics as the primary locus of failure and motivates the behavioural analysis in the following section.
Joint inspection of policy entropy, policy sharpness, and task performance reveals a consistent causal ordering: internal policy degeneration occurs before reward collapse. The agent enters an over-confident, low-uncertainty regime while performance is still improving, and only subsequently does task performance deteriorate irreversibly. This demonstrates that reward collapse is a downstream symptom rather than the root cause of learning failure, and that training-time explainability is required to identify failure mechanisms before they become externally observable.
VI-I Behavioural Lock-In and Dominant Action Dynamics
Figure 15: Temporal evolution of the policy’s dominant action (argmax of episode-mean action probabilities), illustrating early exploratory instability followed by persistent behavioural lock-in.
Figure 15 grounds the internal failure mechanisms identified in Sections VI-G to VI-H in observable agent behaviour. Whereas prior analyses revealed collapse in exploration and confidence dynamics within the policy, this figure demonstrates how those internal pathologies manifest externally as persistent action repetition.
To characterise behavioural evolution over training, we compute episode-wise state-mean policy logits and record the dominant action for each episode. During early training (Episodes 1–18), the dominant action fluctuates rapidly, reflecting sustained exploration and the absence of stable behavioural commitment. Around Episodes 18–19, the policy enters a transient fixation phase, repeatedly selecting Action 6 for several consecutive episodes (Episodes 18–27). This marks a sharp reduction in behavioural diversity and coincides with the initial entropy collapse and confidence amplification identified in Section VI-G. At Episode 28, a second and terminal transition occurs. The policy switches to Action 11 (connect(GIT)) and remains persistently committed for the remainder of training (Episodes 28–50). From this point onward, behavioural adaptation ceases entirely, marking a practically irreversible stage of policy degeneration. This persistent lock-in aligns with the joint explainability analysis in Section VI-H, where internal policy collapse precedes and precipitates task-level performance failure.
This behavioural freezing is independently corroborated by PPO action-distribution snapshots (e.g., Figure 12, CC500 Episode 40), which show the contraction of an initially broad action distribution into a single dominant choice. Together, these results demonstrate that internal policy collapse translates directly into concrete behavioural failure, completing the explainability chain from internal dynamics to external outcomes.
Explainability signals do not merely describe internal learning states; they predict concrete behavioural outcomes. Once policy entropy collapses and decision confidence saturates, the agent transitions into persistent action repetition from which recovery does not occur within training. Behavioural lock-in therefore serves as the observable endpoint of internal policy failure, validating the operational relevance of training-time explainability.
VII Discussion
Key Insights. Across all experiments, three consistent insights emerge. First, training-time signals expose phase transitions, instability, and behavioural collapse that remain invisible under reward-only evaluation, demonstrating that cumulative reward is an unreliable indicator of learning health in adversarial cyber environments. Second, commonly used reinforcement learning enhancements, most notably Prioritised Experience Replay (PER) and entropy-driven exploration in policy-gradient methods, can exhibit pathological behaviour in sparse-reward, partially observable settings. While PER is often assumed to accelerate learning or promote earlier lateral movement, our analysis shows that it instead amplifies transient, high-TD-error transitions, introducing replay bias that distorts value propagation and destabilises strategy formation in large environments. Similar sensitivity arises in policy-gradient methods, where aggressive entropy decay can induce premature determinism rather than sustained exploration. More broadly, our analysis suggests that richer or denser reward signals do not prevent premature escalation, but can instead obscure early over-commitment by maintaining stable reward trajectories despite collapsing exploration. Third, these failure modes emerge early during training, often well before any degradation in task-level performance is observable. Together, these findings establish training-time explainability as a critical diagnostic capability for understanding, debugging, and tuning autonomous cyber attacker agents.